# On an Optimal Stopping Problem with a Discontinuous Reward

### v2

@@ 3.2. Surrender Right - Non-Trivial Case
Clearly, an optimal stopping time for the problem with the continuous reward function is $\widetilde{\tau}_{0}^{x}=0$ since $h(0,x)=\mathbb{E}[e^{-rT}\max(G,F_{T}^{x})]=v(0,x)$; whereas $\bar{\tau}_{0}^{x}=T$ as per (18).
From Example 3.3, it is clear that the three stopping times $\widetilde{\tau}_{t}^{x}$, $\tau_{t}^{x}$ and $\bar{\tau}_{t}^{x}$ are not equal as soon as for some $(t,x)\in[0,T]\times\mathbb{R}_{+}$, $v(t,x)=h(t,x)$ and $h(t,x)>xg(t,x)$. The results below provide a condition on $C$ and $g$ under which the three stopping times are equal.
We first define the surrender region for the problems with the discontinuous and the continuous reward functions, respectively, by
(20) $\widetilde{\mathcal{S}}=\left\{(t,x)\in[0,T)\times\mathbb{R}_{+}|v(t,x)=g(t,x)x\vee h(t,x)\right\}.$
Lemma 3.5.
If the value function $v$ is continuous on $[0,T]\times\mathbb{R}_{+}$ and $\mathcal{S}=\widetilde{\mathcal{S}}$, then $\widetilde{\tau}_{t}^{x}=\tau_{t}^{x}=\bar{\tau}_{t}^{x}$ almost surely for all $0\leq t\leq T$.
The proof of Lemma 3.5 is based on the following result.
Lemma 3.6.
For $(t,x)\in\mathcal{S}$, $g(t,x)x\geq h(t,x)$.
Proof.
For $(t,x)\in[0,T)\times\mathbb{R}_{+}$, we know that
$v(t,x)=\sup_{\tau\in\mathcal{T}_{t,T}}\mathbb{E}\left[e^{-r(\tau-t)}\varphi(\tau,F_{\tau}^{t,x})\right]\geq\mathbb{E}\left[e^{-r(T-t)}\varphi(T,F_{T}^{t,x})\right]=h(t,x),$
and that on $\mathcal{S}$, $v(t,x)=\varphi(t,x)=g(t,x)x$. The result follows. ∎
We can now prove Lemma 3.5.
Lemma 3.5.
The stopping times $\tau^{x}_{t}$ and $\widetilde{\tau}^{x}_{t}$ can be written as
$\begin{array}[]{cc}\tau^{x}_{t}&=\inf\{t\leq s\leq T|(s,F_{s}^{t,x})\in\mathcal{S}\}\wedge T,\\ \widetilde{\tau}^{x}_{t}&=\inf\{t\leq s\leq T|(s,F_{s}^{t,x})\in\widetilde{\mathcal{S}}\}\wedge T.\end{array}$
Thus, $\tau^{x}_{t}(\omega)=\widetilde{\tau}^{x}_{t}(\omega)$ for all $\omega\in\Omega$, since $\mathcal{S}=\widetilde{\mathcal{S}}$ by assumption. From the continuity of $v$, it follows that $\mathcal{S}$ and $\widetilde{\mathcal{S}}$ are closed sets and since $\mathcal{S}=\widetilde{\mathcal{S}}$, $g(t,x)x\geq h(t,x)$ for any $(t,x)\in\widetilde{\mathcal{S}}$ (by Lemma 3.6). It follows from the definition of $\bar{\tau}_{t}^{x}$ that $\bar{\tau}_{t}^{x}=\widetilde{\tau}_{t}^{x}$, a.s. ∎
Remark 3.5.
In Example 3.3, $\mathcal{S}=\emptyset$ and $\widetilde{\mathcal{S}}=[0,T)\times\mathbb{R}_{+}$.
In the next section, we show that under additional assumption on the drift of (4), the value function is continuous and thus satisfies one of the conditions of Lemma 3.5. We also present a condition on the fee and surrender charge functions under which $\mathcal{S}=\widetilde{\mathcal{S}}$. This result is particularly interesting since it provides a condition under which the two optimal stopping problems defined in (7) and (14) are equivalent. That is, the two problems have the same value function (Theorem 3.4), surrender region (Proposition 4.9), and optimal stopping time (Corollary 4.4).
@@ 4. Analytical study of the value function
In Section 4.1 we study the properties of the value function of the original problem defined in (7). In Section 4.2, we establish the relationship between $v$, a free-boundary value problem, and a variational inequality, analogously to Amarican option pricing. To do so, we follow the reasoning of Jacka (1991) and Lamberton (1998) in the context of American put option pricing. This allows us to derive, in Section 4.3, two other representations for the value function: the surrender premium representation which is analogous to the exercise premium representation in the American option pricing terminology, and the continuation premium representation. This representation is new to the literature on VA and American option pricing and allows us to characterize the (non-)emptiness of the surrender region in Section 4.4. Section 4.5 presents a condition on the fee and the surrender charge functions under which the optimal stopping problem with discontinuous reward function defined in (7) is equivalent to the one with the continuous reward function, in (14). Assumption 4.1 concerns the fee function and how it impacts the drift term in the account value process (4), which we denote by ${\mu:[0,T]\times\mathbb{R}_{+}\rightarrow\mathbb{R}}$, with $\mu(t,x)=(r-C(t,x))x$.
@@
    The fee function $C$ is locally Hölder continuous in $t$.
@@
Many results in this section require Assumptions 4.1 and 4.2 to hold. However, note that as soon as Assumption 4.2 holds, the drift term in (4) is globally Lipschitz in $x$, so Assumption 4.1(i) is automatically satisfied.
Assumption 4.3.
The surrender function $g$ has bounded first-order derivatives on $[0,T]\times\mathbb{R}_{+}$.
We define the continuation region by
(21) $\mathcal{C}=\{(t,x)\in[0,\,T)\times\mathbb{R}_{+}:v(t,x)>\varphi(t,x)\}.$
It follows from (19) that $\mathcal{C}\cup\mathcal{S}=[0,T)\times\mathbb{R}_{+}$, since $v(t,x)\geq\varphi(t,x)$ for all $(t,x)\in[0,T]\times\mathbb{R}_{+}$.
@@ 4.1. Elementary Properties of the Value Function
If Assumption 4.1 (i) holds, the value function $v$ is continuous on $[0,T]\times\mathbb{R}_{+}$.
@@
The continuity of the reward function implies the continuity of the value function (see Krylov (1980), Theorem 3.1.8). Thus, the assertion follows from the continuous reward representation of the value function in Theorem 3.4 and Remark 3.4. ∎
@@
The first assertion follows from Doob’s maximal inequality. Assertions (ii) and (iii) follow easily from (7).
@@
If the fee function is bounded from below by a constant $k>0$, then the upper bound for $v$ can be made sharper using the maximal inequality for geometric Brownian motion of Graversen and Peskir (1998), Theorem 5.1. More precisely, we have that
$v(t,x)\leq G+x\left[1+\frac{\sigma^{2}}{2k}-\frac{\sigma^{2}}{2k}\exp\left(-\frac{(\sigma^{2}+2k)^{2}}{2\sigma^{2}}(T-t)-1\right)\right].$
@@
Under Assumption 4.2, the function $x\mapsto v(t,\,x)$ is non-decreasing and convex for all $t\in[0,T]$.
@@
The price of American options is usually non-increasing in time. This will always be the case if the underlying asset price process and the reward function are time-homogeneous, which is true for call and put options under the Black-Scholes model. However, in our setting, the reward function is time-dependent and may be increasing in $t$ (because of the surrender charge function $g$). Therefore, the value function is not necessarily monotonic in time.
We now examine the smoothness of the value function. The next two results are inspired by Lamberton (1998), Proposition 10.3.5 in the context of the American put option price in the Black-Scholes model. We show that the results still hold in our setting despite significant differences in the reward function, namely its time-dependence and the discontinuity at $t=T$.
@@
    Let Assumptions 4.2 and 4.3 hold. For every $x\in\mathbb{R}_{+}$, there exist constants $C_{1},\,C_{2}>0$ (which may depend on $x$) satisfying,
    $|v(t,x)-v(s,x)|\leq C_{1}|\sqrt{T-t}-\sqrt{T-s}|+C_{2}|t-s|,$
@@
Assertion (i) follows from the Lipschitz property of $\varphi$, whereas (ii) follows from Proposition 10.3.5 (1) of Lamberton (1998) and the Lipschitz property of the exponential function on a disk.
@@
In Proposition 4.1 (i), the assumption that the surrender charge function only depends on time (Assumption 4.2) can be relaxed to more general state-dependent surrender charge functions as long as $x\mapsto xg(t,x)$ remains Lipschitz continuous.
@@ 4.2. Free Boundary Value Problem and Variational Inequality
Using Theorem 3.4 and the previously established properties of the value function, we can apply classical results from optimal stopping theory to establish the relationship between $v$, a free-boundary value problem, and a variational inequality. From the definition of the surrender and continuation regions in (19) and (21), we can deduce from the continuity of the value function (Theorem 4.1) and the continuity of the reward function on the interval $[0,T)\times\mathbb{R}_{+}$ that the continuation region is open and the surrender region is closed. The exercise boundary (or surrender boundary) is the boundary $\partial\mathcal{C}$ of $\mathcal{C}$. Henceforth, the $t$-section of the surrender (resp. continuation) region is denoted by $\mathcal{S}_{t}$ (resp. $\mathcal{C}_{t}$). That is, for $t\in[0,\,T)$,
For $0\leq t\leq T$, we define the second-order differential operator $\mathcal{L}_{t}$ by
@@
Note that $L$ is the term on the right-hand side of (9). When the fee and the surrender charge functions only depend on time (that is, when Assumption 4.2 holds), $L$ becomes
@@
Let Assumption 4.1 hold. The value function $v$ is the unique solution to the boundary value problem
@@
In particular, the function $v\in C^{1,2}$ on $\mathcal{C}$.
Once the continuity of the value function has been established (Theorem 4.1), the optimal stopping time has been shown to exist (Corollary 3.1), and the martingale property of the Snell envelope of the discounted reward process has been proven (see Theorem 3.3 (ii)), the connection between partial differential equations and optimal stopping problems can be established assuming enough regularity of the coefficients in (4). A sufficient condition for the results of Friedman (1964) to hold is that the drift term in (4) is locally Hölder continuous, which is why Assumption 4.1 is needed in Theorem 4.2. The proof makes use of standard partial differential equation (PDE) results in solving the Dirichlet (or the first initial-boundary value) problem, see for instance Friedman (1964), Theorem 3.4.9. For details of the proof in the context of the American put option pricing in the Black-Scholes setting, see Jacka (1991), Proposition 2.6 or Karatzas and Shreve (1998), Theorem 2.7.7, among others. For more general results, one can also refer to Peskir and Shiryaev (2006), Sections 3.7.1 and 4.8.2 or Jacka and Lynn (1992), Proposition 3.1. Assumption 4.1 is not sufficient to provide much information on the shape of the surrender region. However, we show in the next section (Proposition 4.8) that further conditions can be imposed on the fee and the surrender charge functions in order to express the surrender region explicitly in terms of the optimal surrender boundary $b$. Under these assumptions, the free-boundary value problem can be stated more explicitly.
@@
Let Assumptions 4.1 and 4.2 hold. If $L(t)<0$ for all $t\in[0,T)$, then, the value function $v$ is the unique solution to the boundary value problem
@@
with $b(t):=\inf(\mathcal{S}_{t})$. In particular, the function $v$ is $C^{1,2}$ on $\mathcal{C}$.
Assuming $L(t)<0$ for all $t\in[0,T)$ in Corollary 4.1 ensures that $\mathcal{S}\neq\emptyset$ (see Proposition 4.6). The proof of this corollary is a direct consequence of Theorem 4.2 and the results of the next section in Proposition 4.8.
@@
For each $t\in[0,T)$, the condition $L(t)<0$ is equivalent to $\frac{\,\mathrm{d}\ln g(t)}{\,\mathrm{d}t}<c(t)$.
@@
[…]
Next, we discuss the link between the value function and a variational inequality. In Proposition 4.2 below, we establish a result that is a direct consequence of the process $J=\{J_{t}\}_{0\leq t\leq T}$, with $J_{t}:=e^{-rt}v(t,\,F_{t})$, being the Snell envelope of the discounted reward process $Z$; thus, $J$ is a supermartingale by Theorem 3.3 (i).
@@
Let Assumption 4.2 hold. The value function $v$ defined in (7) satisfies
@@
for all $(t,x)\in(0,T)\times\mathbb{R}_{+}$ (with all derivatives in the sense of distribution).
This proposition is analogous to Proposition 10.3.7 of Lamberton (1998) in the context of American put option pricing. For an introduction to weak derivatives and distribution theory see Evans (2010), Chapter 5, and Rudin (1991), Chapter 6.
@@
The idea of the proof of Proposition 4.2 rests on the assumption that the value function is regular enough to apply Itô’s lemma to $J$ (which is usually not the case since $v$ may not be smooth enough across the boundary, see Remark 4.6). $\mathcal{L}_{t}v+v_{t}-rv$, being the drift term of $J$, must be non-positive since $J$ is a supermartingale. In the context of American put option pricing, a rigorous proof relying on distribution theory is given in Lamberton (1998), Proposition 10.3.7. Their arguments can easily be adapted to our problem when the fee is time-dependent. The result also holds for more general surrender functions depending on the sub-account value. Thus, in Proposition 4.2, Assumption 4.2 can be relaxed to more general surrender charge functions.
Remark 4.8.
Other authors have obtained results similar to Proposition 4.2 in more general settings, see for instance Bensoussan and Lions (1982) Section 3.4.9 or Jaillet et al. (1990), Theorem 3.1 and 3.2. However, the results of Bensoussan and Lions (1982) require the function ($(t,x)\mapsto xg(t,x)$) to be bounded, and the ones of Jaillet et al. (1990) consider a continuous and time-independent reward function. Applying the results of Jaillet et al. (1990) to the continuous reward representation (14) by adding a second dimension (time) to the underlying process, as is often done for time-dependent reward functions, is not directly possible because some conditions with the time derivative are not satisfied by the continuous reward function, the first-order derivative of $h$ being unbounded. Establishing the results of Proposition 4.2 in a more general market model is thus left as future research.
Once (25) has been established, we can show that the value function solves a variational inequality when Assumptions 4.1 and 4.2 hold.
@@
with the terminal condition $v(T,x)=\max(G,x)$.
The proof is a direct consequence of Theorem 4.2 and Proposition 4.2.
@@
Let Assumptions 4.2, and 4.3 hold. The partial derivatives $\partial v/\partial x$, $\partial v/\partial t$ and $\partial^{2}v/\partial x^{2}$ are locally bounded, and $\partial v/\partial x$ is continuous on $[0,T)\times\mathbb{R}_{+}$.
@@
Since the value function is Lipschitz (see Proposition 4.1), its first derivatives (in the sense of distribution) in $t$ and in $x$ are locally bounded. Using the convexity of $x\mapsto v(t,x)$ (Lemma 4.2) and the local boundedness of the first order derivatives and (25), we can show that the second order derive in $x$ is also locally bounded (see for instance Lamberton (1998), Theorem 10.3.8 or Jaillet et al. (1990), Theorem 3.6 for details). Following the argument of Lamberton (1998), Corollary 10.3.10 (or Jaillet et al. (1990), Corollary 3.7), we can conclude that the function $\frac{\partial v}{\partial x}$ is continuous on $[0,T)\times\mathbb{R}_{+}$, see Ladyz̆enskaja et al. (1968), Chapter 2, Lemma 3.1. This is also known as the smooth fit conditions in the American option terminology. ∎

### v3

@@ 3.2. Surrender Right - Non-Trivial Case
Clearly, an optimal stopping time for the problem with the continuous reward function is $\widetilde{\tau}_{0}^{x}=0$ since $h(0,x)=\mathbb{E}[e^{-rT}\max(G,F_{T}^{x})]=v(0,x)$; whereas $\bar{\tau}_{0}^{x}=T$ as per (15), since $v(0,x)>g(0,x)x$.
In the next section, we present a condition on the fee and surrender charge functions under which $\widetilde{\tau}_{t}^{x}=\tau_{t}^{x}$ a.s. This result is particularly interesting since it provides a condition under which the two optimal stopping problems defined in (7) and (13) are equivalent. That is, the two problems have the same value function (Theorem 3.4), surrender region (Proposition 4.8), and optimal stopping time (Corollary 4.3).
@@ 4. Analytical study of the value function
The representation of the value function in terms of a continuous reward function simplifies the analysis of its regularity, which is presented in Section 4.1. However, additional work is required to assess its smoothness, which we discuss in Proposition 4.1 below. In Section 4.2, we establish the relationship between $v$, a free-boundary value problem, and a variational inequality. To do so, we use the continuity of the value function and follow the reasoning of Jacka (1991) and Lamberton (1998) in the context of American put option pricing. This allows us to derive, in Section 4.3, two other representations for the value function: the surrender premium representation which is analogous to the exercise premium representation in the American option pricing terminology, and the continuation premium representation. This representation is new to the literature on VA and American option pricing and allows us to characterize the (non-)emptiness of the surrender region in Section 4.4. Section 4.5 presents a condition on the fee and the surrender charge functions under which the optimal stopping problem with discontinuous reward function defined in (7) is equivalent to the one with the continuous reward function, in (13).
Assumptions 4.1 and 4.2 presented next concern the fee and the surrender function and are needed for several of the results of this section to hold. In Assumption 4.1 below, ${\mu:[0,T]\times\mathbb{R}_{+}\rightarrow\mathbb{R}}$ denotes the drift term of the account value process (4), that is, $\mu(t,x)=(r-C(t,x))x$.
@@
    For all $x$, the fee function $C(\cdot,x)$ is locally Hölder continuous.
@@
As soon as Assumption 4.2 holds, the drift term in (4) is globally Lipschitz in $x$, so Assumption 4.1(i) is automatically satisfied. However, we still write the two assumptions separately, since for some of the results in this section, the assumptions can be relaxed and only one of the two assumptions is needed.
For the rest of this paper, we assume that Assumptions 4.1 and 4.2 hold. This restricts the applicability of some of our results. For example, Assumptions 4.1(i) and 4.2, preclude the use of the state-dependent fee function of Bernard et al. (2014a). Not all of our results require both assumptions to hold; when this is the case, it will be indicated in a remark after the result.
@@ 4.1. Elementary Properties of the Value Function
An important result for the analysis of the value function is its continuity, which is easily obtained from its alternative representation in terms of a continuous reward function.
@@
The value function $v$ is continuous on $[0,T]\times\mathbb{R}_{+}$.
@@
The continuity of the reward function implies the continuity of the value function (see Krylov (1980), Theorem 3.1.8, which requires Lipschitz continuity of the drift term in (4), hence the need for Assumption 4.1(i)). Thus, the assertion follows from the continuous reward representation of the value function in Theorem 3.4 and Remark 3.3. ∎
Remark 4.2.
Theorem 4.1 only requires Assumption 4.1 (i) to hold. In particular, the value function is continuous even when the fee and surrender functions depend on both time and the account value, as long as the drift term of (4) is globally Lipschitz in $x$.
@@
The first assertion follows from optional sampling since $\{e^{-r(u-t)}F_{u}^{t,x}\}_{u=t}^{T}$ is a supermartingale. Assertions (ii) and (iii) follow easily from (7).
@@
The properties given in Lemma 4.1 follow from the definition of the contract and hold in a more general setting. They do not require Assumptions 4.1 and 4.2 to be satisfied.
@@
The function $x\mapsto v(t,\,x)$ is non-decreasing and convex for all $t\in[0,T]$.
@@
Only Assumption 4.2 is necessary for Lemma 4.2 to hold.
Remark 4.5.
The value function in the American option pricing problem is usually non-increasing in time. This will always be the case if the underlying asset price process and the reward function are time-homogeneous, which is true for call and put options under the Black-Scholes model. However, in our setting, the reward function is time-dependent and may be increasing in $t$ (because of the surrender charge function $g$). Thus, the value function is not necessarily monotone in time.
Proposition 4.1 below concerns the smoothness of the value function and is inspired by Lamberton (1998), Proposition 10.3.5 in the context of the American put under Black-Scholes. Our proof, which can be found in the Appendix, presents non-trivial differences compared to the American put case, which are due to the time-dependence of the value function.
@@
    For every $x\in\mathbb{R}_{+}$, there exist constants $C_{1},\,C_{2}>0$ (which may depend on $x$) satisfying,
    $|v(t,x)-v(s,x)|\leq C_{1}|t-s|+C_{2}|\sqrt{T-t}-\sqrt{T-s}|,$
@@
Proposition 4.1 holds in a more general setting. Part (i) requires that the fee function depends only on time and that the surrender charge function is such that $x\mapsto xg(t,x)$ remains Lipschitz continuous. Part (ii) requires that Assumption 4.2 is satisfied.
@@ 4.2. Free Boundary Value Problem and Variational Inequality
While there exist significant differences between our problem and the well-studied pricing of an American option, the continuous reward representation of Theorem 3.4 and the previously established continuity of the value function allows us to apply classical results from optimal stopping theory. In this section, we use this strategy to establish the relationship between the value function $v$ associated with the original optimal stopping problem 19, a free-boundary value problem, and a variational inequality. Thus, the proofs of the results presented in this section are kept short; the reader is referred to the literature for more details.
The value function is supported on $[0,T]\times\mathbb{R}_{+}$, which is separated into the surrender region $\mathcal{S}$ and the continuation region $\mathcal{C}$, defined by
(17) $\mathcal{C}=\{(t,x)\in[0,\,T)\times\mathbb{R}_{+}:v(t,x)>\varphi(t,x)\},$
respectively. It follows that $\mathcal{C}\cup\mathcal{S}=[0,T)\times\mathbb{R}_{+}$, since $v(t,x)\geq\varphi(t,x)$ for all $(t,x)\in[0,T]\times\mathbb{R}_{+}$. Furthermore, since $v$ and $\varphi$ are continuous on $[0,T)\times\mathbb{R}_{+}$ (see Theorem 4.1) , we have that the continuation region is open and the surrender region is closed. The exercise (or surrender) boundary is the boundary $\partial\mathcal{C}$ of $\mathcal{C}$.
To express the value function as the solution of a free-boundary problem, we define the second-order differential operator $\mathcal{L}_{t}$, for $0\leq t\leq T$, by
@@
Note that $L$ is the term on the right-hand side of (9). We give this general definition of $L$ since many of our results apply in a more general setting. However, under Assumption 4.2, which is assumed to hold throughout Section 4, $L$ becomes
@@
The next results confirm that $v(t,x)$ solves a free-boundary problem.
@@
The value function $v$ is the unique solution to the boundary value problem
@@
That is, any solution $v$ to (19) can be expressed by (7). In particular, the function $v\in C^{1,2}$ on $\mathcal{C}$.
Remark 4.7.
Theorem 4.2 only requires Assumption 4.1 to hold. In particular, even when the fee and surrender charge functions depend on time and the value of the account, it may still be possible to express the value function as a solution to a problem of the form (19).
Once the continuity of the value function has been established (Theorem 4.1), the optimal stopping time has been shown to exist (Corollary 3.1), and the martingale property of the Snell envelope of the discounted reward process has been proven (see Theorem 3.3 (ii), the connection between partial differential equations and optimal stopping problems can be established assuming enough regularity of the coefficients in (4). A sufficient condition for the results of Friedman (1964) to hold is that the drift term in (4) is locally Hölder continuous, which is why Assumption 4.1 is needed in Theorem 4.2. The proof makes use of standard partial differential equation (PDE) results in solving the Dirichlet (or the first initial-boundary value) problem, see for instance Friedman (1964), Theorem 3.4.9.
The main difficulty in establishing the free-boundary value problem resides in proving the continuity of the value function (Theorem 4.1). Once this is done, the free-boundary value problem follows from standard arguments. Specifically, the continuity of the value function implies that the continuation region is open. Then, the Dirichlet problem (or the first initial-boundary value problem) can be posed in an open subset (typically a ball or a rectangle) of the continuation region. Since the boundary of the defined open subsets is sufficiently regular, standard PDE results guarantee that the Dirichlet problem admits a unique solution. Probabilistic arguments are then used to identify this solution as the value function.
For details of the proof in the context of the American put in the Black-Scholes setting, see Jacka (1991), Proposition 2.6 or Karatzas and Shreve (1998), Theorem 2.7.7, among others. For more general results, see Peskir and Shiryaev (2006), Sections 3.7.1 and 4.8.2 or Jacka and Lynn (1992), Proposition 3.1.
Assumption 4.1 does not give much information on the shape of the surrender region. However, we show in Section 4.4 (Theorem 4.3) that under further conditions on the fee and the surrender charge functions, the surrender region can be expressed in terms of the optimal surrender boundary $b:[0,T]\mapsto\mathbb{R}_{+}$. Then, the free-boundary value problem can be stated more explicitly.
@@
If $L(t)<0$ for all $t\in[0,T)$, the value function $v$ is the unique solution to the boundary value problem
@@
with $b(t):=\inf(\mathcal{S}_{t})$. That is, any solution $v$ to (19) can be expressed by (7). In particular, the function $v$ is $C^{1,2}$ on $\mathcal{C}$.
Assuming $L(t)<0$ for all $t\in[0,T)$ in Corollary 4.1 ensures that $\mathcal{S}\neq\emptyset$ (see Proposition 4.6). This corollary is a direct consequence of Theorem 4.2 and Theorem 4.3(i) of Section 4.4.
@@
For each $t\in[0,T)$, $L(t)<0$ is equivalent to $\frac{\,\mathrm{d}\ln g(t)}{\,\mathrm{d}t}<c(t)$.
@@
When $L(t)<0$ for all $t\in[0,T]$, we have from Corollary 4.1 that $v\in C^{1,2}$ on the continuation region, that is, when $(t,x)\in[0,T]\times(0,b(t))$. On the other hand, when $(t,x)\in[0,T]\times[b(t),\infty)$, we know that $v(t,x)=g(t)x$, so $v\in C^{1,2}$ on $\text{int}(\mathcal{S})$. What remains to assess is the smoothness of the value function across the boundary (i.e. when $x=b(t)$). When $L(t)\geq 0$ for all $t\in[0,T]$, an optimal stopping time for (7) is $T$, as per Proposition 3.1. Hence, $v(t,x)=h(t,x)$, the present value of the maturity benefit (5). The regularity of the value function then follows from the Feynman-Kac representation (see Karatzas and Shreve (1991), Theorem 5.7.6), which states that, under some regularity assumptions on the coefficients of (4), $h$ is the unique solution to a Cauchy problem, and such a solution is in $C^{1,2}([0,T)\times\mathbb{R}_{+})$.
Next, we discuss the link between the value function and a variational inequality. In Proposition 4.2 below, we establish a result that is a direct consequence of the process $J=\{J_{t}\}_{0\leq t\leq T}$, with $J_{t}:=e^{-rt}v(t,\,F_{t})$, being the Snell envelope of the discounted reward process $Z$, and thus a supermartingale by Theorem 3.3 (i).
@@
The value function $v$ defined in (7) satisfies
@@
for all $(t,x)\in(0,T)\times\mathbb{R}_{+}$, with all derivatives in the sense of distribution.
Proposition 4.2 is analogous to Proposition 10.3.7 of Lamberton (1998) in the context of the American put. For an introduction to weak derivatives and distribution theory see Evans (2010), Chapter 5, and Rudin (1991), Chapter 6.
The idea of the proof of Proposition 4.2 rests on the assumption that the value function is regular enough to apply Itô’s lemma to $J$ (which is usually not the case since $v$ may not be smooth enough across the boundary, see Remark 4.9). $\mathcal{L}_{t}v+v_{t}-rv$, being the drift term of $J$, must be non-positive since $J$ is a supermartingale. In the context of American put option pricing, a rigorous proof relying on distribution theory is given in Lamberton (1998), Proposition 10.3.7. Their arguments can easily be adapted to our problem, as long as the fee function does not depend on the account value.
@@
Proposition 4.2 holds in a more general setting, as it only requires Assumption 4.2 to be satisfied. In fact, its proof only needs the fee function to be independent of the account value; the condition on the surrender charge function in Assumption 4.2 can be relaxed.
From Theorem 4.2 and Proposition 4.2, we obtain that the value function $v$ solves a variational inequality.
@@
with terminal condition $v(T,x)=\max(G,x)$ and all derivatives in the sense of distribution.
Remark 4.11.
Proposition 4.3 holds as long as Assumption 4.1 is satisfied and the fee function does not depend on $x$.
@@
The partial derivatives (in the sense of distribution) $\partial v/\partial x$, $\partial v/\partial t$ and $\partial^{2}v/\partial x^{2}$ are locally bounded, and $\partial v/\partial x$ is continuous on $[0,T)\times\mathbb{R}_{+}$.
@@
From Proposition 4.1, we have that the first order derivatives of the value function (in the sense of distribution) in $t$ and in $x$ are locally bounded. Using the convexity of $x\mapsto v(t,x)$ (Lemma 4.2), the local boundedness of the first order derivatives and (21), we can show that the second order derivative in $x$ is also locally bounded (see for instance Lamberton (1998), Theorem 10.3.8 or Jaillet et al. (1990), Theorem 3.6 for details). This last result is also known as the smooth fit condition in the American option terminology.
The continuity of $\frac{\partial v}{\partial x}$ on $[0,T)\times\mathbb{R}_{+}$ is obtained using the first part of the statement; any function $u(t,x)$ defined on $\mathbb{R}\times\mathbb{R}$ has a continuous first partial derivative in $x$ if $\partial u/\partial x$, $\partial u/\partial t$ and $\partial^{2}u/\partial x^{2}$ are bounded, see Ladyz̆enskaja et al. (1968), Chapter 2, Lemma 3.1. Indeed, Lipschitz-continuity in $x$ comes from the boundedness of $\partial^{2}U/\partial x^{2}$, and the boundedness of both $\partial u/\partial t$ and $\partial^{2}u/\partial x^{2}$ are used to show continuity in $t$ (see the proof of Corollary 10.3.10 of Lamberton (1998) for a sketch of the proof).